OpenStudy (anonymous):

Applying the Mean Value Theorem with a = 1, b=4 to the function in the figures below leads to c=2. what is the equation of the tangent line at 2?

OpenStudy (dumbcow):

you can get tangent line from using point-slope form \[y - y_1 = m(x-x_1)\] x1 = 2 y1 = f(2) m = dy/dx = f'(2)

OpenStudy (anonymous):